OpenStudy (anonymous):

A tunnel with a parabolic arch is 12m wide. If the height of the arch 4m from the left edge is 6m, can a truck that is 5m tall and 3.5 m wide pass through the tunnel? HELP!

OpenStudy (anonymous):

I kind of have the answer but I don't know what to do next..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

Think about it if 4m away it its 6m tall then 6m away its going to grow. Every two meters twords the middle it grows three meters up

OpenStudy (anonymous):

y = a(x-r)(x-s) = a (x-12)(x-0) 6 = a(4-12)(4-0) 6 = a(-8)(4) 6 = -32a -- ---- -32 -32 6 = a - -- 32 - 3 -- = a 16

OpenStudy (anonymous):

but I need to prove my justification using quadratic equations.

OpenStudy (anonymous):

Ok so you know it can fit width wise. So we need to find height now

OpenStudy (anonymous):

2nd: y = -3/16 x(x-12) 5 = -3/16x(x-12) 16*5 = -3/16*16x(x-12) 80 = -3x^2 + 36x 3x^2 -36x + 80 = 0 a = 3, b = -36, c = 80 x = -(-36) +/- √(-36)^2 - 4(3)(80) --------------------------- 2(3)

OpenStudy (anonymous):

= 36 +/- √336 ------------ 6 x = 36 + √336 ----------- 6 = 9.1 or x= 36 - √336 ---------- 6 = 2.9

OpenStudy (anonymous):

that kinda makes sense. o me this is not a well written question considering its given the car can fit but Ill try and help

OpenStudy (anonymous):

oh, thank you Jacob. & the answer key says: f(x) = -3/16x (x-12) Yes, because at a height of 5 m the bridge is 6.11 m wide. How is it 6.11 m wide? o.o

OpenStudy (anonymous):

I forgot what my teacher told me eep >_<"

OpenStudy (anonymous):

Oh ok. Think of this as a parabola for the y value of 5 the x value is 6.11. Th bridge up to that point is 6.11m wide

OpenStudy (anonymous):

Does that kinda make sense?

OpenStudy (anonymous):

no not really.. D:

OpenStudy (anonymous):

ohhh, wait yes I get that the x value will be 6.11 but how? o.o

OpenStudy (anonymous):

try plugging 5 as the y value to one of your equatons

OpenStudy (anonymous):

I tried adding the zeros and divide by 2 but it's 10.31 not 6.11 o.o

OpenStudy (anonymous):

What you need to do is find the equation of the parabola with some of the given points

OpenStudy (anonymous):

then plug in 5

OpenStudy (anonymous):

That's what I did in the beginning and I'm still confused as to how they got 6.11...

OpenStudy (anonymous):

I'm sorry. I dont know the answer I can ask my teacher tommorow if you want he can help.